Melissa is making scale models of 2 buildings. The Empire State Building is 1250 feet tall. Melissa's model of the Em pire State Building is 2 meters tall. The seattle space Needle is 605 feet tall. If Melissa uses the same scale, which proportion can be used to find h the height in meters of her model of the Space Needle?
Answer by checkley71(8403)

2/1250=X/605 NOW CROSS MULTIPLY.
1250X=2*605
1250X=1210
X=1210/1250
X=.968 METERS FOR THE SCALE MODEL FOR THE 605 FOOT BUILDING.